OpenStudy (anonymous):

2x - 4 + 4x + 10 = 9x - 9 What's the value of x? Please help! Thanks!

OpenStudy (anonymous):

hey, I saw you posted a similar problem earlier. It's solved the exact same way. Start by combining like terms.

OpenStudy (baby_bear69):

2x-4+4x+10= 9x-9 6x-6 = 9x-9 +6 +6 6x = 9x-3 -9x -9x -3x= -3 --- -- -3 -3 x= 1

OpenStudy (baby_bear69):

That don't seem right tho, what are your choices?`

OpenStudy (anonymous):

it's not...when you combined like terms on the left -4 + 10 equals +6, not -6